# SkillMimic: Learning Reusable Basketball Skills from Demonstrations

### v1

@@ Abstract.
Mastering basketball skills such as diverse layups and dribbling involves complex interactions with the ball and requires real-time adjustments. Traditional reinforcement learning methods for interaction skills rely on labor-intensive, manually designed rewards that do not generalize well across different skills. Inspired by how humans learn from demonstrations, we propose SkillMimic, a data-driven approach that mimics both human and ball motions to learn a wide variety of basketball skills. SkillMimic employs a unified configuration to learn diverse skills from human-ball motion datasets, with skill diversity and generalization improving as the dataset grows. This approach allows training a single policy to learn multiple skills, enabling smooth skill switching even if these switches are not present in the reference dataset. The skills acquired by SkillMimic can be easily reused by a high-level controller to accomplish complex basketball tasks. To evaluate our approach, we introduce two basketball datasets: one estimated through monocular RGB videos and the other using advanced motion capture equipment, collectively containing about 35 minutes of diverse basketball skills. Experiments show that our method can effectively learn various basketball skills included in the dataset with a unified configuration, including various styles of dribbling, layups, and shooting. Furthermore, by training a high-level controller to reuse the acquired skills, we can achieve complex basketball tasks such as layup scoring, which involves dribbling toward the basket, timing the dribble and layup to score, retrieving the rebound, and repeating the process. Project page and video demonstrations: https://ingrid789.github.io/SkillMimic/
character animation, human-object interaction, reinforcement learning, humanoid robot, basketball
submissionid: 1234ccs: Computing methodologies Procedural animationccs: Computing methodologies Control methods
@@
Figure 1. We propose a novel approach that enables physically simulated humanoids to learn a variety of basketball skills from human-object demonstrations, such as shooting (blue), retrieving (red), and turnaround layup (yellow). Once acquired, these skills can be reused and combined to accomplish complex tasks, such as continuous scoring (green), which involves dribbling toward the basket, timing the dribble and layup to score, retrieving the rebound, and repeating.
1. Introduction
Basketball, as a collection of highly dynamic, complex, diverse, and precise object-interaction skills, exemplifies the pinnacle of human athleticism and control capabilities. Professional basketball players spend extensive time practicing fundamental skills such as shooting and dribbling, which they flexibly utilize to score during games. These basketball skills typically involve not only isolated body movements but also interactions with the ball, requiring real-time adjustments based on the state of the object to achieve precise control. Current state-of-the-art methods apply reinforcement learning to mimic isolated body movements, combining these motion priors with manually designed skill rewards to learn interaction skills, such as striking pillars (Peng et al., 2022), playing tennis (Zhang et al., 2023b), climbing ropes (Bae et al., 2023), and carrying boxes (Hassan et al., 2023). However, designing these skill rewards is labor-intensive and difficult to generalize across different skills. Given the complexity and diversity of basketball skills, creating skill rewards for each basketball skill demands significant engineering effort. Consequently, existing methods fail to provide a unified learning framework for a single policy to learn diverse basketball skills, let alone the challenge of reusing and composing these skills to accomplish more complex tasks, such as continuous scoring.
How can a simulated humanoid learn a wide variety of reusable basketball skills in a simple and scalable manner? We can draw inspiration from the way humans learn basketball skills from demonstrations. Most basketball players, even without formal coaching, can master various skills by watching basketball videos and practicing diligently. During this learning process, individuals adjust their body movements to align with both the reference body motions and the reference ball movements they observe. This paradigm of learning skills from demonstrations allows a basketball player to acquire a set of reusable basketball skills and flexibly combine them for further purposes.
@@ 1. Introduction
Inspired by these observations, we propose a data-driven method called SkillMimic, which simultaneously mimics human and object motions and their contacts to learn interaction skills. SkillMimic can learn various basketball skills purely from human-ball motion datasets, such as diverse styles of shooting, layups, and dribbling skills, and even a robust pickup skill that enables picking up balls in random locations and motions. Notably, SkillMimic uses the exact same configuration to learn different skills, with identical hyperparameters. This allows us to train a single policy to learn multiple skills and achieve smooth skill switching, even if these switches are not present in the reference dataset. By training a high-level policy to reuse these learned skills, we can accomplish challenging high-level tasks such as scoring, which requires the humanoid to composite diverse basketball skills to score accurately.
To evaluate our method, we introduce two basketball datasets. The first dataset, BallPlay-V, uses a purely visual approach, employing neural networks to estimate the motion of players and basketballs from RGB videos, and includes eight basic basketball skills. The second dataset, BallPlay-M, utilizes advanced optical motion capture equipment to record the movements of both players and balls, encompassing approximately 35 minutes of comprehensive basketball skills. Experiments on BallPlay-V demonstrate that our method exhibits robust performance against data inaccuracies. In BallPlay-M, our system, SkillMimic, successfully learns various reusable basketball skills within a single policy using a unified configuration. By additionally training a high-level controller to flexibly compose the skills acquired by SkillMimic, we can efficiently complete complex tasks such as directional dribbling and layup scoring. Compared to previous methods, our approach offers significant advantages in simplicity and efficiency in learning basketball skills and complex basketball tasks.
@@
  • SkillMimic: A data-driven paradigm for learning reusable interaction skills, capable of learning diverse basketball skills in a unified manner. It supports joint learning and smooth switching of skills, with skill diversity and generalization improving as the dataset grows.
@@
  • Contact Graph: We propose a simple and general contact modeling method that applies to diverse skills, called the contact graph. A Contact Graph Reward (CGR) is designed to enable precise contact imitation, which proves to be critical for learning precise interaction skills.
@@
  • Unified skill imitation reward: We propose a set of important designs that form a unified reward configuration for imitation learning of various interaction skills.
@@
  • A hierarchical solution for learning complex basketball tasks: We propose training a high-level controller to flexibly compose the skills acquired by SkillMimic to accomplish challenging high-level tasks.
@@
  • BallPlay Datasets: We introduce two basketball datasets to facilitate research on basketball skill learning.
2. Related Work
2.1. Human Motion Imitation and Skill Learning
@@
Mimicking human motion for robot control is an efficient method of learning humanoid skills. DeepMimic (Peng et al., 2018), a pioneer in this field, uses imitation learning to perform a variety of highly dynamic skills, including backflips, cartwheels, and running. The introduction of Generative Adversarial Imitation Learning (GAIL) (Ho and Ermon, 2016) into humanoid imitation learning by AMP (Peng et al., 2021) lessens constraints on data alignment, thus enhancing its versatility. ASE (Peng et al., 2022) further amplifies motion diversity in GAIL training, incorporates a pre-trained low-level policy to acquire locomotion skills, and deploys a high-level policy to repurpose these locomotion skills for specific tasks, such as striking pillars at random locations. This methodology is subsequently expanded in numerous efforts to augment condition and text control (Juravsky et al., 2022; Tessler et al., 2023; Dou et al., 2023; Zhu et al., 2023; Sun et al., 2023; Ren et al., 2024b; Pan et al., 2024), as well as to design automated rewards (Cui et al., 2024). This collection of studies employs imitation learning to broaden the scope of data-driven skill acquisition, enabling the control and reuse of these learned skills for high-level tasks. However, a limitation of these studies is their exclusive focus on learning isolated locomotion skills through imitation. This requires additional pipeline and reward design for learning interaction skills, even for relatively simple ones, such as moving boxes (Hassan et al., 2023) or striking pillars (Peng et al., 2022). In contrast, our approach achieves unified learning of diverse basketball interaction skills without the need for skill-specific reward design and hyperparameter tuning.
2.2. Generating Human-Object Interactions
@@
Owing to the complex dynamics involved in diverse interactions, the generation of interaction for simulated humanoids has long posed challenges in the fields of graphics and robotics. Early studies often relied on manually designed control structures. For example, Hodgins et al. (Hodgins et al., 1995) implement a variety of sports interactions, including running, bicycling, and vaulting, through the use of manually designed state machine controllers. Yin et al. (Yin et al., 2007) use finite-state machines to generate a large variety of gaits. Coros et al. (Coros et al., 2010) utilize an inverted pendulum model to control foot placement, thereby achieving a variety of gait controls. With the advancement of machine learning, deep reinforcement learning (Arulkumaran et al., 2017) has been widely employed in character control, giving rise to an array of methods based on policy network control. These methods have been applied to a range of challenging activities, such as basketball (Liu and Hodgins, 2018), skateboarding (Liu and Hodgins, 2017), and cycling (Tan et al., 2014), etc. The integration of imitation learning further simplifies the learning of interactions, and this data-driven learning paradigm significantly lessens the need for intricate expert design. Zhang et al. (Zhang et al., 2023b) propose a framework to learn diverse tennis skills from broadcast videos. Bae et al. (Bae et al., 2023) incorporate multiple part-wise motion priors for diverse whole-body interactions, such as climbing ropes and weightlifting. Similarly, Braun et al. (Braun et al., 2023) develop a framework for synthesizing whole-body motions for dexterous grasps.
A closely related work is (Zhang et al., 2023a), where an interaction graph (Ho et al., 2010) that represents the relative motions is proposed to learn the simulated multi-character interaction. However, its kinematic-only rewards fail to measure correct contacts and are unstable in learning diverse interactions. Instead, our proposed contact-aware imitation rewards prioritize contact learning and significantly improve the success rate. Additionally, a policy trained by (Zhang et al., 2023a) is limited to strictly mimicking a single reference clip during testing, which is insufficient for it to be considered a skill. In contrast, our method can train on multiple clips and learn generalizable and controllable skills beyond mere imitation. Another highly related work is (Braun et al., 2023), a framework for physically plausible whole-body grasping. Similar to previous kinematics-based grasping synthesis methods (Wu et al., 2022; Ghosh et al., 2023), they design skill-specific contact rewards as guidance for learning grasp skills. However, the method contains multiple networks and training rounds with pre-trained modules dedicated to the grasp skills. Beyond grasp, our method covers diverse basketball interaction skills, within a simple unified framework, and without designing skill-specific rewards.
In addition to physics-based methods, a multitude of interaction generation approaches rooted in kinematics are flourishing (Li et al., 2023b, a; Xu et al., 2023, 2024; Jiang et al., 2024, 2023; Starke et al., 2020, 2021). These methods typically use networks to predict HOI states without incorporating physical simulations. Consequently, the generated interactions often lack physical authenticity and demand large volumes of data.
3. Preliminaries on Reinforcement Learning
Our method is based on reinforcement learning, where the agent interacts with the environment according to a policy to maximize the expected return. The basic task is formulated as a Markov Decision Process (MDP) defined by the tuple =$\mathcal{M}=\{\boldsymbol{s},\boldsymbol{a},\boldsymbol{f},r,\gamma\}$ of states, actions, transition dynamics, reward function, and discount factor. At each time step $t$, the agent takes the system states $\boldsymbol{s}_{t}$ as input and outputs an action $\boldsymbol{a}_{t}$ by sampling the policy distribution $\boldsymbol{\pi}(\boldsymbol{a}_{t}|\boldsymbol{s}_{t})$. The action $\boldsymbol{a}_{t}$ will result in a new state $\boldsymbol{s}_{t+1}$ through physical simulation $\boldsymbol{f}(\boldsymbol{s}_{t+1}|\boldsymbol{a}_{t},\boldsymbol{s}_{t})$. Then a reward $r_{t}=r(\boldsymbol{s}_{t},\boldsymbol{a}_{t},\boldsymbol{s}_{t+1})$ is calculated, and the goal is to learn a policy that maximizes the expected return $\mathcal{R}(\boldsymbol{\pi})=\mathbb{E}_{p_{\boldsymbol{\pi}}(\boldsymbol{\tau})}\left[\sum_{t=0}^{T-1}\gamma^{t}r_{t}\right]$, where $\boldsymbol{\tau}=\{\boldsymbol{s}_{0},\boldsymbol{a}_{0},r_{0},...,\boldsymbol{s}_{T-1},\boldsymbol{a}_{T-1},r_{T-1},\boldsymbol{s}_{T}\}$ represents the trajectory, $p_{\boldsymbol{\pi}}(\boldsymbol{\tau})$ is the probability density function of the trajectory. $T$ denotes the time horizon, and $\gamma\in[0,1]$ is a discount factor. To enable skill control, we optionally provide the policy with extra conditions $\boldsymbol{c}$, yielding a variant policy $\boldsymbol{\pi}(\boldsymbol{a}_{t}|\boldsymbol{s}_{t},\boldsymbol{c})$.
4. BallPlay Dataset
4.1. BallPlay-V
The BallPlay-V dataset applies a monocular annotation solution to estimate the high-quality human SMPL-X (Pavlakos et al., 2019) parameters and object translations from RGB videos. However, annotating these videos with high-speed and dynamic movements and complex interactions in the 3D camera coordinate is quite challenging. Inspired by the whole-body annotation pipeline of Motion-X (Lin et al., 2023), our automatic annotation additionally introduces depth estimation (Bhat et al., 2023), semantic segmentation (Ren et al., 2024a), to obtain high-quality whole-body human motions and ball motions. The BallPlay-V dataset contains eight basketball skills, including back dribble, cross leg, hold, fingertip spin, pass, backspin, cross, and rebound.
4.2. BallPlay-M
Although BallPlay-V can acquire HOI data conveniently from RGB videos, its accuracy is limited due to significant depth estimation errors. Additionally, it struggles with occlusion issues, making it difficult to capture complex layup and dribbling data. To achieve more comprehensive and accurate basketball data, we create the BallPlay-M dataset using optical motion capture equipment. This dataset comprises 35 minutes of diverse human-basketball interactions, including 39 dribble clips, 131 pickup clips, 8 layup clips, 13 shot clips, 10 pass clips, 16 getup clips, and 21 miscellaneous clips. During the capture process, optical markers are attached to both the player and the basketball to track body and ball movements. The player also wears gloves equipped with inertia sensors to estimate finger movements. Consequently, the player is parameterized as a skeleton with 52 joints (156 DOFs). We calculate and record the root rotation, root translation, joint positions, and joint rotations sequentially. The ball is parameterized as a sphere, with its rotation and center position recorded. All data are captured at 120 fps.
@@ 4.2. BallPlay-M
Figure 2. Concept of SkillMimic. We define a skill as a specific set of Human-Object Interaction (HOI) state transitions that align with the intended skill semantics. These state transitions can be derived from capturing real-world skills into HOI motion clips. If a simulated humanoid can manipulate objects such that the resulting HOI state transitions closely match those of the reference, we consider the humanoid to have successfully learned the skill.
5. SkillMimic
In this section, we introduce a unified data-driven framework for simulated humanoids to learn various reusable basketball skills. We begin by introducing how we define skills with Human-Object Interaction (HOI) data (Sec. 5.1). Next, we explain how to learn diverse skills by imitating HOI data (Sec. 5.2) and discuss the designing of a unified HOI imitation reward (Sec. 5.3). Finally, we demonstrate how the learned skills can be reused to perform complex tasks (Sec. 5.4).
5.1. Defining Skills Using Data
We propose a data-driven approach to defining skills as collections of Human-Object Interaction (HOI) state transitions that align with the intended skill semantics. For instance, the skill of ”picking up balls from various positions” can be encapsulated by various HOI motion clips, each capturing a unique instance of the action. If we consider each frame of HOI data as a state, these clips collectively form a collection of state transitions that represent the skill.
Unlike defining skills using explicit rules, our approach uses data to represent skills. This makes it easier to define skills that are difficult to describe with rules, such as various basketball skills. In addition, by collecting richer skill data samples, the skill coverage can be more completely represented. For example, a pickup skill defined by 100 different HOI motion clips will be more representative than one defined by a single clip.
5.2. Learning Skills by Imitation
Considering skills defined by a set of reference HOI state transitions, if a humanoid can manipulate objects such that its HOI state transitions closely resemble those of the reference, we consider the humanoid to have successfully learned the skill. Based on this concept, we propose a method for learning interaction skills by imitating HOI state transitions, which we call SkillMimic. Fig. 2 illustrates the concept of SkillMimic. Unlike previous approaches (Peng et al., 2018, 2021, 2022; Zhang et al., 2023b) that require manually designed skill rewards for each interaction skill, SkillMimic is fully data-driven, skill-agnostic, and scalable, making it capable of learning a wide range of basketball skills within a unified solution. Fig. 6 shows diverse basketball skills acquired using SkillMimic. Fig. 8 and Fig. 9 illustrate how the performance of the pickup skill improves as the amount of reference data increases.
Thanks to the exploratory nature of deep reinforcement learning, the skills learned through SkillMimic can quickly switch from minor out-of-domain states to learned state distributions, leading to skill generalization and zero-shot skill switching.
5.2.1. Training.
Fig. 3 (b) shows the training pipeline of SkillMimic. Given an HOI dataset with diverse skill-labeled clips (Fig. 3 (a)), SkillMimic trains by randomly selecting a frame from a clip corresponding to a chosen skill to initialize the humanoid and object states. The state $\boldsymbol{s}_{t}$ (Sec. 5.2.2) and skill label $\boldsymbol{c}_{j}$ are input into the policy, which predicts actions (Sec. 5.2.3) that are then simulated to produce the next state $\boldsymbol{s}_{t+1}$. A unified HOI imitation reward (Sec. 5.3) is designed to measure the consistency between the simulated state transitions and the reference transitions. When the simulation triggers specific conditions, such as falling or reaching the maximum simulation length, the environment resets, and the process repeats. After training converges, the policy enables the humanoid to execute skills that closely mirror the reference demonstrations. Moreover, since it does not rely on reference data during testing, skills can operate continuously with robustness and generalization capabilities.
Our implementation is based on Isaac Gym (Makoviychuk et al., 2021), which supports the parallel execution of numerous environments. To quickly adapt to the diverse states encapsulated in the dataset, we apply the reference state initialization (RSI) (Peng et al., 2018) independently for each environment. Since clips vary in length, the rewards at convergence can differ significantly, potentially causing imbalanced skill learning. To ensure more balanced skill learning, we set a fixed maximum simulation length, making the upper limit of imitation rewards similar across all skills.
@@ 5.2. Learning Skills by Imitation
Figure 3. Our system consists of three parts. (a) First, we capture real-world basketball skills to create a large Human-Object Interaction (HOI) motion dataset. (b) Second, we train a skill policy to learn interaction skills by imitating the corresponding HOI data. Specifically, the policy takes as input the HOI state $\boldsymbol{s}_{t}$ and skill label $\boldsymbol{c}_{j}$ and predicts the action $\boldsymbol{a}_{t}$. The new state $\boldsymbol{s}_{t+1}$ is calculated by the simulator. A unified HOI imitation reward is designed to imitate diverse HOI state transitions. (c) The third part involves training a High-Level Controller (HLC) to reuse the learned skills for complex tasks. The HLC takes as input $\boldsymbol{s}_{t}$ and extra task observations $\boldsymbol{h}_{t}$, e.g., the basket position, and predicts the skill label $\boldsymbol{c}_{t}$ to drives a pre-trained skill policy.
5.2.2. Observation.

### v2

# SkillMimic: Learning Basketball Interaction Skills from Demonstrations

@@ Abstract
Traditional reinforcement learning methods for human-object interaction (HOI) rely on labor-intensive, manually designed skill rewards that do not generalize well across different interactions. We introduce SkillMimic, a unified data-driven framework that fundamentally changes how agents learn interaction skills by eliminating the need for skill-specific rewards. Our key insight is that a unified HOI imitation reward can effectively capture the essence of diverse interaction patterns from HOI datasets. This enables SkillMimic to learn a single policy that not only masters multiple interaction skills but also facilitates skill transitions, with both diversity and generalization improving as the HOI dataset grows. For evaluation, we collect and introduce two basketball datasets containing approximately 35 minutes of diverse basketball skills. Extensive experiments show that SkillMimic successfully masters a wide range of basketball skills including stylistic variations in dribbling, layup, and shooting. Moreover, these learned skills can be effectively composed by a high-level controller to accomplish complex and long-horizon tasks such as consecutive scoring, opening new possibilities for scalable and generalizable interaction skill learning. Project page: https://ingrid789.github.io/SkillMimic/
@@
Figure 1: We propose a novel approach that for the first time enables physically simulated humanoids to learn a variety of basketball interaction skills from Human-Object Interaction (HOI) data, including but not limited to shooting (blue), retrieving (red), and turnaround layup (yellow). Once acquired, these interaction skills can be composed to accomplish complex tasks, such as consecutive scoring (green).
00footnotetext: Equal Contribution. Corresponding author.
1 Introduction
Human-Object Interaction (HOI) represents one of the most fundamental yet challenging capabilities in robotics and character animation [17, 11, 52, 24, 30, 51, 13]. Among various HOI scenarios, basketball serves as an exemplary testbed that demands mastery of complex, dynamic, and precise interactions. Despite significant progress, existing physics-based character animation methods predominantly focus on imitation learning of locomotion skills [36, 37, 38, 9, 47] through Reinforcement Learning (RL), leaving a significant gap for HOI imitation. While some methods can learn interaction skills, such as striking pillars [38], playing tennis [55], climbing ropes [2], and carrying boxes [16], they typically involve handcrafted reward functions for each different interaction skill. However, designing skill-specific rewards is labor-intensive and fundamentally limits generalization. This limitation becomes particularly acute in basketball, where the complexity and diversity of skills would demand extraordinary engineering effort for reward design. Consequently, existing methods [36, 37, 38, 9, 47, 55, 2, 16] fail to provide a unified learning framework for a single policy to learn diverse interaction skills, let alone achieving complex long-term tasks such as consecutive basketball scoring.
How can a simulated humanoid learn a wide variety of interaction skills in a simple and scalable manner? Consider how humans naturally acquire complex skills: amateur basketball players often develop their expertise by watching and imitating game footage, even without formal coaching. This self-directed learning process involves synchronizing their body movements with both the observed player motions and ball trajectories through practice.
@@ 1 Introduction
Inspired by these observations, we propose a data-driven method called SkillMimic, which mimics HOI through RL to learn interaction skills. SkillMimic can learn various interaction skills purely from HOI datasets, such as diverse styles of basketball shooting, layup, and dribbling skills, and even a robust pickup skill that enables picking up balls in random locations and motions. Notably, SkillMimic uses the exact same configuration to learn different interaction skills, with identical hyperparameters. This allows us to train a single Interaction Skill (IS) policy to learn multiple interaction skills and achieve smooth skill switching. By training a high-level policy to reuse these learned interaction skills, we can accomplish challenging high-level tasks such as scoring, which requires the humanoid to composite diverse basketball interaction skills to score accurately.
To advance basketball skill learning, we introduce two HOI datasets: BallPlay-V, which captures player and ball motion from RGB videos across eight basic skills, and BallPlay-M, which uses optical motion capture to record 35 minutes of comprehensive basketball interactions. Using these datasets, SkillMimic successfully learns diverse basketball skills through a unified IS policy. With an additional high-level controller for skill composition, our system can execute complex tasks like consecutive scoring. Our approach demonstrates superior efficiency in learning both interaction skills and multi-phase, long-horizon tasks compared to existing methods, opening new possibilities for scalable and generalizable interaction skill learning.
@@
  • SkillMimic: a data-driven paradigm for learning diverse humanoid interaction skills through RL. It supports joint learning and switching of diverse basketball interaction skills, with skill diversity and generalization improving as the demonstration data grows, showcasing the scalability.
@@
  • Contact graph: We propose a simple and general contact modeling method that applies to diverse skills, called the contact graph. A contact graph reward is designed to enable precise contact imitation, which proves to be critical for learning precise interaction skills.
@@
  • Unified HOI imitation reward: We propose a set of important designs that form a unified reward configuration for imitation learning of various interaction skills.
@@
  • A hierarchical solution for reusing interaction skills: We propose training a high-level controller to reuse the interaction skills acquired by SkillMimic to accomplish challenging high-level tasks, which for the first time achieves continuous basketball scoring.
@@
  • BallPlay Datasets: We introduce two basketball datasets to facilitate research on interaction skill learning.
2 Related Work
Locomotion Imitation Learning.
@@ 2 Related Work
Mimicking human motion for robot control is an efficient method of learning humanoid skills. DeepMimic [36], a pioneer in this field, uses imitation learning to perform a variety of highly dynamic locomotion skills. The introduction of Generative Adversarial Imitation Learning (GAIL) [19] into humanoid imitation learning by AMP [37] lessens constraints on data alignment, thus enhancing its versatility. ASE [38] further amplifies motion diversity in GAIL training, incorporates a pre-trained low-level policy to acquire locomotion skills, and deploys a high-level policy to repurpose these locomotion skills for specific tasks, such as striking pillars. These methodologies are subsequently expanded in numerous efforts to augment condition and text control [23, 47, 9, 58, 44, 39, 34, 31, 49], as well as to design automated rewards [8]. However, a limitation of these studies is their exclusive focus on learning isolated locomotion skills through imitation. This requires additional pipeline and reward design for learning interaction skills, even for relatively simple ones, such as moving boxes [16] or striking pillars [38]. In contrast, our approach enables unified learning of diverse interaction skills, eliminating the need for skill-specific reward engineering.
Learning Human-Object Interaction (HOI).
@@
Learning HOI in simulation has long posed challenges in the fields of graphics and robotics [25, 24, 52, 53, 22, 21, 42, 43, 50, 15]. Early studies often relied on manually designed control structures. For example, Hodgins et al. [20] implement a variety of sports interactions through the use of manually designed state machine controllers. Yin et al. [54] use finite-state machines to generate a large variety of gaits. Coros et al. [7] utilize an inverted pendulum model to control foot placement. Recently, deep reinforcement learning [1] has been widely employed in character control, giving rise to an array of methods based on policy network control. These methods have been applied to a range of challenging activities, such as basketball [28], skateboarding [27], and cycling [46], etc. The use of imitation rewards further lessens the reward design for locomotion styles. Zhang et al. [55] propose a framework to learn diverse tennis skills from broadcast videos. Bae et al. [2] incorporate multiple part-wise motion priors for diverse whole-body interactions. Similarly, Braun et al. [5] develop a framework for synthesizing whole-body grasps.
Applying imitation learning for HOI is intuitive, however, effectively learning diverse interaction skills in a unified framework remains unresolved. Naively extending locomotion imitation methods [36, 37] to HOI imitation often yields unstable and unnatural results, as these methods do not address the unbalanced rewards inherent in HOI, nor do they properly account for the crucial aspects of relative motion and contact in HOI. Modeling interaction through interaction graphs is a well-established approach [18]. While this method has proven useful for HOI imitation [57], it remains limited by considering only kinematic relationships, failing to ensure proper contact and becoming unstable in learning diverse scenarios. In contrast, our contact graph imitation reward explicitly prioritize physical contact learning, significantly improving the success rate of HOI imitation and enabling unified HOI imitation of diverse complex interaction skills for the first time.
This work builds upon and significantly extends ideas initially explored in our early technical report [48].
@@
Figure 2: Concept of SkillMimic. We define an interaction skill as a set of Human-Object Interaction (HOI) state transitions that align with the intended skill semantics. These state transitions can be derived from captured HOI motion clips. If a simulated humanoid can manipulate objects such that the resulting HOI state transitions closely match those of the reference, we consider the humanoid to have successfully learned the interaction skill.
@@
Figure 3: Our system consists of three parts. (a) First, we capture real-world basketball skills to create a large Human-Object Interaction (HOI) motion dataset. (b) Second, we train an Interaction Skill (IS) policy to learn interaction skills by imitating the corresponding HOI data through reinforcement learning. Specifically, the IS policy takes as input the HOI state $\boldsymbol{s}_{t}$ and skill label $\boldsymbol{c}_{j}$ and predicts the action $\boldsymbol{a}_{t}$. The new state $\boldsymbol{s}_{t+1}$ is calculated by the simulator. A unified HOI imitation reward is designed to imitate diverse HOI state transitions. (c) The third part involves training a High-Level Controller (HLC) to reuse the learned interaction skills for complex tasks. The HLC takes as input $\boldsymbol{s}_{t}$ and extra task observations $\boldsymbol{h}_{t}$, e.g., the basket position, and predicts the skill label $\boldsymbol{c}_{t}$ to drive a pre-trained IS policy.
3 SkillMimic
In this section, we introduce a unified data-driven framework for simulated humanoids to learn various basketball interaction skills via simulation and reinforcement learning (RL). We begin by introducing how we define interaction skills with Human-Object Interaction (HOI) data (Sec. 3.2). Next, we explain how to learn diverse interaction skills by imitating HOI data (Sec. 3.3) and discuss the designing of a unified HOI imitation reward (Sec. 3.4). Finally, we demonstrate how the learned interaction skills can be reused to perform high-level tasks (Sec. 3.5).
3.1 The BallPlay Dataset
To address the scarcity of basketball HOI data and facilitate research on interaction skill learning, we introduce two datasets: BallPlay-V, based on monocular vision estimation, containing eight clips of different basketball skills; and BallPlay-M, based on multi-view optical motion capture systems, containing 35 minutes of diverse basketball interactions. We provide more details in the appendix.
3.2 Defining Interaction Skills Using HOI Data
We propose a data-driven approach to defining interaction skills as collections of Human-Object Interaction (HOI) state transitions that align with the intended skill semantics. For instance, the interaction skill of “picking up balls from various positions” can be encapsulated by various HOI motion clips, each capturing a unique instance of the action. These clips collectively form a collection of HOI state transitions that represent the interaction skill.
3.3 Learning Interaction Skills by HOI Imitation
Considering an interaction skill defined by a set of reference HOI state transitions, if a humanoid can manipulate objects such that its HOI state transitions closely resemble the reference, we consider the humanoid to have successfully learned the interaction skill. Based on this concept, we propose a method for learning interaction skills by imitating HOI state transitions through RL, which we call SkillMimic. Fig. 2 illustrates the concept of SkillMimic.
In contrast to Behavior Clone [4, 12, 6] methods which rely on precise state-action data, SkillMimic learns actions through RL using state-only trajectories, demonstrating better tolerance to data noise and higher data efficiency. Unlike previous RL methods [37, 38, 55] that require manually designed rewards for each interaction skill, SkillMimic is fully data-driven, skill-agnostic, and scalable, making it capable of learning a wide range of basketball skills within a unified solution. Fig. 6 shows diverse basketball skills acquired using SkillMimic. Fig. 7 illustrates how the performance of the pickup skill improves as the reference data increases.
3.3.1 Training Pipeline
Fig. 3 (b) shows the training pipeline of SkillMimic. Given an HOI dataset with diverse skill-labeled clips (Fig. 3 (a)), SkillMimic trains an Interaction Skill (IS) policy by randomly selecting a frame from a clip corresponding to a chosen skill to initialize the humanoid and object states. The state $\boldsymbol{s}_{t}$ (Sec. 3.3.2) and skill label $\boldsymbol{c}_{j}$ are input into the IS policy, which predicts actions (Sec. 3.3.3) that are then simulated to produce the next state $\boldsymbol{s}_{t+1}$. A unified HOI imitation reward (Sec. 3.4) is designed to measure the consistency between the simulated state transitions and the reference transitions. When the humanoid falls or reaches the maximum simulation length, the environment resets, and the process repeats. After training converges, the IS policy enables the humanoid to execute interaction skills that closely mirror the reference demonstrations.
3.3.2 HOI Observation